# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v2

@@ 5.1 Experiment Settings
90.5 $\pm$4.3 89.7
@@
89.9
$\pm$1.4
Average  77.5 84.7 83.8 84.1 85.5  75.1 75.8 80.2 80.2 83.5
5.2 Main Results
Results for Gym Tasks. The experimental results on D4RL Gym tasks are summarized in Table 1. Our proposed BiTrajDiff demonstrates consistent superiority across all dataset configurations. Specifically, it achieves notable performance gains in all of the environments under both TD3BC and IQL algorithms, while simultaneously maintaining high stability. Moreover, BiTrajDiff consistently outperforms existing DA methods across most scenarios. In terms of final average performance, BiTrajDiff achieves gains of 10.3% on IQL and 11.2% on TD3BC, substantially exceeding improvements reported by alternative methods like DiffStitch. These consistent gains across both IQL and TD3BC backbones underscore the effectiveness and generality of our BiTrajDiff approach.
5.3 Visulization
Figure 1: Visualization of the comparative dynamic error and L2 distance metrics for BiTrajDiff-generated trajectories versus forward and backward diffusion-generated trajectories, all evaluated with a consistent horizon.
To evaluate the accuracy and diversity robustness of the BiTrajDiff framework, we compare its generated trajectories with those produced by single-direction forward and backward diffusion models. Following [30], we employ two quantitative metrics to assess a generated trajectory $\hat{\tau}$: 1) Dynamic Error: $\mathcal{E}_{\text{Dyn}}(\hat{\tau})=\sum_{t}\|\hat{s}_{t+1}-s_{t+1}\|_{2}$, which measures trajectory accuracy by summing the mean squared error between the predicted next states $\hat{s}_{t+1}$ and the ground truth $s_{t+1}$; 2) L2 Distance: $\mathcal{E}_{\text{L2D}}(\hat{\tau})=\min_{\tau\in\mathcal{D}}\sum_{t}\|\hat{s}_{t}-s_{t}\|_{2}$, defined as the minimal L2 distance between the generated trajectory and any trajectory of equal length in the dataset $\mathcal{D}$, capturing trajectory diversity. As shown in Figure 1, while forward and backward diffusion models achieve low $\mathcal{E}_{\text{Dyn}}$, they exhibit limited $\mathcal{E}_{\text{L2D}}$, indicating that the generated trajectories closely resemble those in the original dataset and thus offer limited performance gains. In contrast, trajectories generated by BiTrajDiff demonstrate significantly broader marginal distributions in $\mathcal{E}_{\text{L2D}}$ while maintaining competitive $\mathcal{E}_{\text{Dyn}}$, reflecting enhanced behavioral diversity without sacrificing dynamic consistency. Consequently, BiTrajDiff facilitates greater performance gains with the augmented data. These results highlight the ability of BiTrajDiff to generate behaviorally diverse trajectories while preserving accuracy, effectively balancing exploration and reliability in offline trajectory generation.
6 Conclusion
In this paper, we present BiTrajDiff, a novel diffusion-based framework that enhances offline RL by enabling bidirectional trajectory generation from shared intermediate states. Unlike previous single-direction augmentation methods that only model forward future trajectories from observed states, BiTrajDiff simultaneously generates both forward and backward trajectories. This dual-generation strategy allows BiTrajDiff to not only imagine plausible futures but also reconstruct potential historical transitions, thereby creating novel trajectories that globally connect states, including those unreachable in the original dataset. This leads to significantly diverse and effective trajectory augmentation, improving offline RL performance beyond the capability of the forward-only generation schemes of other DA frameworks. Extensive experiments on the D4RL benchmark also demonstrate that BiTrajDiff significantly outperforms existing data augmentation baselines across diverse offline RL algorithms. Future work will explore bidirectional generation for offline RL under multi-task settings with broader behavioral patterns.

### v3

@@ Appendix D Experimental results with the direction of diffusion generation
102.8 $\pm$0.6   64.4
$\pm$8.9
83.5
$\pm$12.9
83.8
$\pm$9.7
85.0 $\pm$10.3
Average  79.1 78.9 82.1 85.5  75.1 78.3 82.7 85.7
Table S5: Test returns of our proposed BiTrajDiff with augmented single- and bi-directional diffusion generated trajectories on CQL and DT algorithms under D4RL locomotion tasks.
Source Task  CQL (kumar2020conservative)  DT (chen2021decision)
  Base Forward Backward Ours  Base Forward Backward Ours
medium halfcheetah    47.4
$\pm$0.2
47.5
$\pm$0.2
48.9 $\pm$0.4 48.7
$\pm$0.3
42.7
$\pm$0.5
42.7
$\pm$0.2
43.0
$\pm$0.3
43.3 $\pm$0.2
walker2d    82.6
$\pm$1.0
80.9
$\pm$2.7
83.8
$\pm$0.7
84.2 $\pm$0.5   69.6
$\pm$11.0
74.6
$\pm$3.6
76.1
$\pm$3.5
77.8 $\pm$4.0
hopper    68.8
$\pm$4.1
71.6
$\pm$14.9
75.2
$\pm$9.5
80.9 $\pm$3.4   56.0
$\pm$1.9
56.6
$\pm$5.6
56.9
$\pm$5.7
57.8 $\pm$2.2
medium expert halfcheetah    87.2
$\pm$4.6
88.0
$\pm$3.9
92.6
$\pm$2.2
94.1 $\pm$1.1   70.8
$\pm$12.3
74.3
$\pm$4.7
64.7
$\pm$4.4
83.5 $\pm$6.3
walker2d    110.2
$\pm$0.6
110.5 $\pm$0.6 109.5
$\pm$0.7
110.4
$\pm$0.4
104.6
$\pm$6.9
105.1
$\pm$5.1
104.9
$\pm$4.8
106.3 $\pm$6.0
hopper    103.5
$\pm$8.1
102.7
$\pm$13.9
102.5
$\pm$7.8
105.7 $\pm$5.8   107.5
$\pm$3.4
108.3
$\pm$3.3
109.8
$\pm$2.4
110.1 $\pm$1.9
medium replay halfcheetah    45.9
$\pm$0.2
45.9
$\pm$0.5
46.2 $\pm$0.4 45.8
$\pm$0.3
38.8
$\pm$2.4
36.4
$\pm$2.9
39.3
$\pm$0.5
39.5 $\pm$0.8
walker2d    82.1
$\pm$1.8
82.9
$\pm$0.6
84.4
$\pm$2.1
84.8 $\pm$1.5   52.1
$\pm$5.5
58.6 $\pm$10.0 54.9
@@
58.2
$\pm$6.1
hopper    95.7
$\pm$2.5
97.9
$\pm$6.1
93.8
$\pm$5.3
100.3 $\pm$1.3   67.3
$\pm$8.9
62.1
$\pm$13.0
68.4
$\pm$17.2
69.6 $\pm$13.1
Average  80.4 80.9 81.9 83.9  67.7 68.7 68.7 71.8
We conduct experiments with different augmented diffusion trajectories generated in single-directional and bi-directional. Table S4 shows the test returns of the IQL and TD3BC algorithms. For CQL and DT algorithms, the quantitative results are reported in Table S5. The algorithms were evaluated on D4RL locomotion tasks with augmented trajectories generated in single-direction and bi-directional. Compared to the base method, the forward method offers only modest gains and is at times even detrimental to performance. Conversely, the backward method generally outperforms the base method across most tasks, achieving substantial gains. Finally, our bi-directional approach, BiTrajDiff, consistently exhibits superior and robust performance. Our BiTrajDiff enables global connectivity between originally unreachable states from the original dataset and thus constructs novel and high-quality augmented trajectories.
Appendix E Experimental results with trajectory filters
Table S6: Test results of BiTrajDiff with data from different variants related to trajectory filters. OOD indicates the OOD filter, and G represents the greedy filter.
Source Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
  Base None
w G,
w/o OOD
w OOD,
w/o G
Ours  Base None
w G,
w/o OOD
w OOD,
w/o G
Ours
medium replay walker2d    73.3
$\pm$7.5
78.5
$\pm$10.2
83.9
$\pm$4.5
80.1
$\pm$7.0
87.6 $\pm$2.2   80.9
$\pm$6.0
77.2
$\pm$24.4
77.5
$\pm$26.9
82.5
$\pm$8.0
89.9 $\pm$1.4
hopper    94.7
$\pm$4.0
93.6
$\pm$12.0
91.6
$\pm$8.2
96.3
$\pm$3.5
102.8 $\pm$0.6   64.4
$\pm$8.9
69.0
$\pm$19.2
75.0
$\pm$17.0
74.6
$\pm$23.6
85.0 $\pm$10.3
We illustrate the necessity of integrating both OOD and greedy filters in our full BiTrajDiff model. While Figure 3d presents the learning curves, Table S6 directly compares the performance on walker2d-medium-replay and hopper-medium-replay tasks under different filter settings. Without any filters, the performance is comparable to, and in some cases inferior to, that of the base method on both the IQL and TD3BC algorithms. When applying the greedy filter in isolation, the performance improvement is noticeable but marred by volatility. In contrast, using the OOD filter alone yields a more stable and substantial performance enhancement. Nevertheless, our full BiTrajDiff model, which integrates both filters, consistently outperforms all other configurations across the tested settings, indicating the necessity of both OOD and greedy filters.
Appendix F Experimental results with augmented data ratio $\sigma$
Figure S1: Performance comparison of BiTrajDiff with different augmented data ratios $\sigma$ in the walker2d-medium-replay task. The best result for each offline RL algorithm is marked as bold.
Ratio     IQL TD3BC CQL
$0\%$     73.3
$\pm$7.5
80.9
$\pm$6.0
82.1
$\pm$1.8
$10\%$     78.7
$\pm$2.4
82.0
$\pm$5.5
84.0
$\pm$3.1
$20\%$     84.2
$\pm$1.3
87.5
$\pm$1.0
83.9
$\pm$2.0
$30\%$     87.6
$\pm$2.2
89.9 $\pm$1.4 84.8
$\pm$1.5
$40\%$     88.7 $\pm$4.5 86.7
$\pm$2.9
84.7
$\pm$2.1
$50\%$     81.6
$\pm$5.6
86.4
$\pm$2.7
86.3 $\pm$2.1
$100\%$     84.5
$\pm$10.2
86.7
$\pm$2.4
83.9
$\pm$3.3
In this section, we conduct experiments to assess the impact of the augmented data ratio $\sigma$. As shown in Table S1, when the ratio $\sigma$ is too low, which is between $0\%\sim 20\%$, performance improvements remain consistently constrained. Conversely, while an augmentation ratio $\sigma$ between $30\%$ and $50\%$ already provides a substantial and stable performance improvement, increasing $\sigma$ to $100\%$ compromises this stability, indicating a decline in the performance. This result demonstrates that both the limited and overabundant synthetic data have detrimental effects on the performance of BiTrajDiff, since the limited data is not enough to provide offline RL algorithms with new behavior patterns, and overabundant data introduces excessive noise, which hinders offline RL algorithms from learning effectively, and an optimal trade-off is achieved with a $30\%$ to $50\%$ augmentation range. As a result, we choose $\sigma=30\%$ as the fixed hyperparameter for all our experiments.
Appendix G Experimental results with $n$-step TD Estimator
Table S7: Test returns of our proposed BiTrajDiff and baselines with varying length of $n$-step TD estimator ($n$) on walker2d-medium-replay task.
Length of n-step TD Estimator($n$)   IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
  base RTDiff diff stitch ours  base RTDiff diff stitch ours
$n=2$  73.3
$\pm$7.5
74.0
$\pm$5.2
74.9
$\pm$8.2
87.6 $\pm$2.2   80.9
$\pm$6.0
83.5
$\pm$7.4
80.5
$\pm$5.2
89.9 $\pm$1.4
$n=5$  71.7
$\pm$12.4
76.4
$\pm$16.6
79.8
$\pm$2.4
86.7 $\pm$0.9   80.3
$\pm$12.0
88.6 $\pm$3.6 87.7
$\pm$4.0
87.3
$\pm$1.3
$n=10$  69.5
$\pm$13.5
73.1
$\pm$12.7
71.4
$\pm$9.9
79.5 $\pm$2.3   81.5
$\pm$7.6
79.6
$\pm$12.7
81.3
$\pm$5.1
87.5 $\pm$3.4
$n=15$  72.7
$\pm$11.4
67.2
$\pm$14.8
69.7
$\pm$6.0
77.9 $\pm$6.2   75.1
$\pm$9.9
80.6
$\pm$13.8
79.1
$\pm$9.7
83.9 $\pm$4.7
Table S8: Test returns of our proposed BiTrajDiff and baselines with varying length of $n$-step TD estimator ($n$) on hopper-medium-replay task.
Length of n-step TD Estimator($n$)   IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
  base RTDiff diff stitch ours  base RTDiff diff stitch ours
$n=2$  94.7
$\pm$4.0
97.7
$\pm$2.7
96.4
$\pm$5.6
102.8 $\pm$0.6   64.4
$\pm$8.9
72.1
$\pm$18.9
71.5
$\pm$21.0
85.0 $\pm$10.3
$n=5$  90.5
$\pm$1.9
94.7
$\pm$5.4
99.9 $\pm$6.3 98.7
$\pm$6.4
67.0
$\pm$96.6
75.3
$\pm$14.8
67.7
$\pm$21.0
89.7 $\pm$12.6
$n=10$  94.5
$\pm$9.8
97.0
$\pm$2.0
93.3
$\pm$9.1
100.0 $\pm$1.3   62.7
$\pm$25.5
66.6
$\pm$15.6
63.3
$\pm$25.7
83.5 $\pm$13.5
$n=15$  82.0
$\pm$26.9
93.3
$\pm$9.1
88.4
$\pm$9.5
99.5 $\pm$6.8   65.0
$\pm$26.0
64.8
$\pm$25.7
39.6
$\pm$17.2
71.3 $\pm$14.3
In this section, we demonstrate the effectiveness of BiTrajDiff in Offline RL algorithms with varying lengths of the $n$-step TD estimators. As shown in Table S7 and Table S8, while RTDiff and diffsitch initially outperform the base method, their performances exhibit a sharp decline, even being surpassed by the base method, as $n$ increases. Conversely, our proposed BiTrajDiff generally achieves the highest scores and consistently maintains its advantage as the $n$-step horizon increases. Meanwhile, as described in section 5.4, the performance of all tested methods degrades with an increasing $n$-step horizon, which is attributed to the escalating variance in value estimation. Moreover, for DA methods, the accumulated errors introduced by the inverse dynamics model should also be considered. In contrast, BiTrajDiff alleviates these issues by synthesizing novel, high-quality long-horizon trajectories, thereby achieving a more stable and consistent performance. This result demonstrates that by stitching forward-future and backward-history trajectories, BiTrajDiff synthesizes higher-quality and more reliable trajectories than existing DA methods.
Appendix H The Use of Large Language Models (LLMs)
In this work, we leverage Large Language Models (LLMs) to enhance and refine textual content. Specifically, LLMs are employed to polish paragraphs, improving their fluency, coherence, grammatical accuracy, and overall readability.